Given a mean of 62 and a standard deviation of 5​, convert the random variables to​ Z-scores, then use a standard normal table below to find the probability.

​Pr[62 ​≤ X​ ≤ 67​] ​=
Answer #1

P(62 < X < 67)

= P((62 - )/< (X - )/< (67 - )/)

= P((62 - 62)/5 < Z < (67 - 62)/5)

= P(0 < Z < 1)

= P(Z < 1) - P(Z < 0)

= 0.8413 - 0.5

= 0.3413